OpenStudy (anonymous):

WILL FAN AND METAL!!!!!! A set of equations is given below: Equation E: a = 5b + 1 Equation F: a = 3b + 4 Which statement describes a step that can be used to find the solution to the set of equations? Equation F can be written as 5b + 1 = 3b + 4. Equation F can be written as b + 1 = 3b + 4. Equation F can be written as a = 3(a − 1) + 4. Equation F can be written as a = 5(a − 4) + 1.

OpenStudy (anonymous):

first person to answer will be fanned and given a metal!

OpenStudy (anonymous):

i also have another question

OpenStudy (anonymous):

look, you can isolate b on equation and substitute on f equation.

OpenStudy (anonymous):

As I remember in algebra, it would be *Equation F can be written as 5b + 1 = 3b + 4.*

OpenStudy (anonymous):

so you'll get b=(a-1)/5 substitute a=3(a-1)/5+4

OpenStudy (anonymous):

i still don't get it